OpenStudy (anonymous):

the expensive ones cost 7$ each, whereas the worthless ones cost only 2$. Monongehela spent 111 and bought 3 more expensive ones than worthless ones. how many of each did she buy

OpenStudy (anonymous):

x= expensive y= worthless x = y+3 (she bought 3 more expensive) 111 = 7x + 2y 111 = 7 (y+3) + 2y 111 = 21y + 7y + 2y 90 = 9 y y = 10 x = 13